# Circuit complexity and functionality: a statistical thermodynamics perspective

### v2

@@ Results
The above intuition concerning fragmentation and polynomial thermalization is further supported by arguments from the theory of computational complexity. According to these arguments, connecting any two circuits of the same functionality and size via a polynomial number of local functionality preserving dynamical moves would imply the equality NP = coNP of two complexity classes, namely NP - decision problems for which polynomial time solutions are not known but for which a YES solution can be checked in polynomial time; and coNP - decision problems for which polynomial time solutions are not known but for which a NO solution can be checked in polynomial time. Even though it is widely believed that NP $\neq$ coNP, rigorously justifying this belief (or its negation) is an open problem in computer science. In order to understand the above claim consider the Circuit Equivalence problem, the problem of deciding whether two (polynomial-sized) circuits are equivalent. Circuit Equivalence is known to be a problem in coNP, since a NO solution can be easily checked by comparing the outputs of the two circuits on a given input. At the same time, if any two circuits could be connected by a polynomial number of dynamical moves, even if difficult to find, this set of moves can be used to verify a YES solution of the Circuit Equivalence problem in polynomial time, placing Circuit Equivalence also in NP. Similarly, a parallel set of arguments can be used to argue that Circuit Inequivalence, i.e., the problem of deciding that two (polynomial-sized) circuits are inequivalent, which is a problem in NP, is also in class coNP since the connectivity of any two circuits via a polynomial number of moves could be used to verify the NO solution of Circuit Inequivalence (i.e., equivalence) in polynomial time. Moreover, it is also well known that Circuit Equivalence and Circuit Inequivalence are among the hardest problems in their respective coNP and NP classes, i.e., they are in classes coNP-complete and NP-complete, respectively. “Completeness” indicates that any problem in coNP or NP can be reduced, respectively, to Circuit Equivalence or Circuit Inequivalence in polynomial time. Given that a polynomial number of dynamical moves placed Circuit Equivalence in NP and Circuit Inequivalence in coNP the conclusions of the above line of argumentation are that: (i) all problems in coNP are in NP (coNP $\subset$ NP); (ii) all problems in NP are in coNP (NP $\subset$ coNP); and thus that (iii) NP = coNP. As already alluded to above, this conclusion contradicts widely accepted beliefs in computational complexity and implies that, in our context, fragmentation is unavoidable, regardless of whether the polynomial sequence of moves connecting pairs of circuits is easy or hard to find.
@@
In summary, the thermodynamic perspective to complexity and functionality of circuits proposed here charters a new line of inquiry. Developing this direction further and addressing quantitatively the questions it raises will require combining the intuition and tools of theoretical physics with mathematical ideas and techniques from the theory of classical and quantum computation. We believe that this work provides a framework that should stimulate new ways of thinking and fruitful interactions at the interface between physics and computer science. In particular, the issue of fragmentation which, to our knowledge, has not been explored by computer scientists, is a topic of much current interest to the physics communities working on classical and quantum dynamics [21, 22, 23, 24, 25]. Conversely, the problem of dynamics of systems with “multiplicative” rather than additive conservation laws (as is the case with the functionality of circuits) defines a new intriguing problem for theoretical physicists interested in classical and quantum dynamics.
@@ Acknowledgments
We are grateful to Alexsey Khudorozhkov and Guilherme Delfino for insightful discussions. This work was supported in part by DOE Grant DE-FG02-06ER46316 (C.C.) and a Grant from the Mass Tech Collaborative Innovation Institute (A.E.R.). R.C., C.C., and A.E.R. also acknowledge the Quantum Convergence Focused Research Program, funded by the Rafik B. Hariri Institute at Boston University.
@@
for which the choice of $\epsilon$ scales with ${\cal N}$ via an infinitesimal constant $\delta$, chosen so that the error $\epsilon=\delta\,{\cal N}$ is small for circuits of a bounded size (e.g., polynomial on the number $n$ of qubits).  [^6] We note that the issues discussed above concerning the need of introducing the “tolerance”, $\epsilon$, in the definition of $\Omega_{\epsilon}$ in Eq. (18) has a direct analogue in physical systems. Specifically, when defining a microcanonical potential for a physical system, $\Omega^{\rm ph}_{\epsilon}(E,N)$, which counts the number of microstates of given energy $E$ and particle number $N$, it is necessary to bracket states of energy between $E-\epsilon/2$ and $E+\epsilon/2$, defining an interval of size $\epsilon$ around the value of $E$. The inequality $\Omega^{\rm ph}_{\epsilon_{1}}(E_{1},N_{1})\;\Omega^{\rm ph}_{\epsilon_{2}}(E_{2},N_{2})\leq\Omega^{\rm ph}_{\epsilon_{1}+\epsilon_{2}}(E_{1}+E_{2},N_{1}+N_{2})$ then follows. Two comments are in order: first, notice that choosing $\epsilon=\delta\,N$ is natural as it keeps the intensive energy $E/N$ within a fixed interval of size $\delta$. Second, the actual choice of the $\epsilon$ is immaterial in statistical mechanics in the thermodynamic limit, because $\epsilon$ corresponds to the thickness of an energy shell in a $N$-dimensional space, and therefore the correction to the entropy from that shell is subextensive. As a result, details on the specific choice of $\epsilon$ (beyond the definitional need to introduce it) are inconsequential.
Eq. (21) then translates into an inequality for $\bar{\Omega}_{\delta}$,
@@
It then follows from Eq. (25) and the inequality Eq. (23) that
@@
B - Model for the dependence of the circuit entropy on the depth and complexity of computation
@@
where $C(n,0)=\binom{n}{n/2}-\binom{n}{n/2+1}$ is a Catalan number that counts the number of balanced non-negative walks, i.e., walks that start and end at $k=0$ after $n$ (even) steps but that can only explore points with $k\geq 0$. By using Pascal’s triangle relations it is then easy to see that the solution to the recursion relation of Eq. (30) is given by
@@
Let us consider the expression Eq. (33) with $p_{+}\gg p_{-},p_{0}$, a limit which is justified because for a generic circuit one expects that, upon addition a new (random) gate the complexity is much more likely to increase than to decrease or even remain unchanged. Also, in this limit $n_{+}\gg n_{-},n_{0}$ and $\frac{n_{+}-n_{-}+1}{n_{+}+1}\to 1$, leading to the following expression for the normalization,
@@
Eq. (40) then reduces to: $\bar{S}(0,{\cal N})\geq{\cal N}\;\log|G|-\frac{1}{2}\;{\cal N}\;\log|G|=\frac{1}{2}\;{\cal N}\;\log|G|$, which recovers the lower bound for the entropy of circuits implementing the identity permutation that we derived in the body of the paper by other means.

### v3

@@ Results
The intuition about fragmentation and polynomial thermalization is supported by computational complexity theory. Connecting two circuits of the same size and functionality by a polynomial number of local, functionality-preserving moves would imply NP = coNP, contradicting widely accepted beliefs of computational complexity theory. To understand this statement, recall that the complexity class NP contains YES decision problems that can be verified in polynomial time, while the class coNP contains problems for which NO solutions can be verified in polynomial time. If two circuits can be connected via a polynomial sequence of moves, Circuit Equivalence, a problem in coNP, is then also in NP because the sequence is itself a witness for the YES decision. Similarly, connecting two circuits through a polynomial sequence of moves implies that Circuit Inequivalence, a problem in NP, is also in coNP. Moreover, it is also well known that Circuit Equivalence and Circuit Inequivalence are among the hardest problems in their respective coNP and NP classes, i.e., they are in classes coNP-complete and NP-complete, respectively. “Completeness” indicates that any problem in coNP or NP can be reduced, respectively, to Circuit Equivalence or Circuit Inequivalence in polynomial time. Given that a polynomial number of moves placed Circuit Equivalence in NP and Circuit Inequivalence in coNP the conclusions of the above line of argumentation is that NP = coNP, implying that fragmentation of the space of ${\cal N}$ gate circuits implementing a permutation $P$ is inevitable, regardless of whether the sequence of moves is easy or hard to find.
@@
In summary, the thermodynamic perspective to complexity and functionality of circuits proposed here charts a new line of inquiry. Developing this direction further and addressing quantitatively the questions it raises will require combining the intuition and tools of theoretical physics with mathematical ideas and techniques from the theory of classical and quantum computation. We believe that this work provides a framework that should stimulate new ways of thinking and fruitful interactions at the interface between physics and computer science. In particular, the issue of fragmentation which, to our knowledge, has not been explored by computer scientists, is a topic of much current interest to the physics communities working on classical and quantum dynamics [26, 27, 28, 29, 30]. Conversely, the problem of dynamics of systems with “multiplicative” rather than additive conservation laws (as is the case with the functionality of circuits) defines a new intriguing problem for theoretical physicists interested in classical and quantum dynamics.
@@ Acknowledgments
We are grateful to Alexsey Khudorozhkov and Guilherme Delfino for insightful discussions. This work was supported in part by NSF Grants GCR-2428487 (A.E.R, C.C., R.C.) and GCR-2428488 (E.R.M), DOE Grant DE-FG02-06ER46316 (C.C.) and a Grant from the Mass Tech Collaborative Innovation Institute (A.E.R.). R.C., C.C., and A.E.R. also acknowledge the Quantum Convergence Focused Research Program, funded by the Rafik B. Hariri Institute at Boston University.
@@
for which the choice of $\epsilon$ scales with ${\cal N}$ via an infinitesimal constant $\delta$, chosen so that the error $\epsilon=\delta\,{\cal N}$ is small for circuits of a bounded size (e.g., polynomial on the number $n$ of qubits).  [^7] We note that the issues discussed above concerning the need of introducing the “tolerance”, $\epsilon$, in the definition of $\Omega_{\epsilon}$ in  (18) has a direct analogue in physical systems. Specifically, when defining a microcanonical potential for a physical system, $\Omega^{\rm ph}_{\epsilon}(E,N)$, which counts the number of microstates of given energy $E$ and particle number $N$, it is necessary to bracket states of energy between $E-\epsilon/2$ and $E+\epsilon/2$, defining an interval of size $\epsilon$ around the value of $E$. The inequality $\Omega^{\rm ph}_{\epsilon_{1}}(E_{1},N_{1})\;\Omega^{\rm ph}_{\epsilon_{2}}(E_{2},N_{2})\leq\Omega^{\rm ph}_{\epsilon_{1}+\epsilon_{2}}(E_{1}+E_{2},N_{1}+N_{2})$ then follows. Two comments are in order: first, notice that choosing $\epsilon=\delta\,N$ is natural as it keeps the intensive energy $E/N$ within a fixed interval of size $\delta$. Second, the actual choice of the $\epsilon$ is immaterial in statistical mechanics in the thermodynamic limit, because $\epsilon$ corresponds to the thickness of an energy shell in a $N$-dimensional space, and therefore the correction to the entropy from that shell is subextensive. As a result, details on the specific choice of $\epsilon$ (beyond the definitional need to introduce it) are inconsequential.
(21) then translates into an inequality for $\bar{\Omega}_{\delta}$,
@@
It then follows from  (25) and the inequality  (23) that
@@
B - Extensivity of circuit entropy at the complexity threshold
Here we argue that the circuit entropy ${\cal S}(P,{\cal K}(P))$ for circuits of minimum length is extensive in the circuit complexity ${\cal K}(P)$. The upper bound ${\cal S}(P,{\cal K}(P))\leq{\cal K}(P)\;\log_{2}|G|$ derived from Eq. (5) of the main text establishes that the circuit entropy is at most extensive in ${\cal K}(P)$. Below we argue that ${\cal S}(P,{\cal K}(P))$ is also bounded below by a term proportional to ${\cal K}(P)$. For concreteness, we use the gate set $G$ comprised of 3-bit permutation gates in $S_{8}$, with $|G|=8!\binom{n}{3}$. We consider circuits for which all inputs are affected by at least one gate, i.e., none of the $n$ bitlines are untouched by the circuit.
We proceed by considering the outputs of each of the ${\cal K}(P)$ gates in the circuit. Each of these outputs are either: (a) connected to the inputs of another gate; or (b) exit unimpeded as output bitlines of the circuit. In case (a) we have an internal link between the two gates. There are $3\times{\cal K}(P)$ outputs originating from the ${\cal K}(P)$ 3-bit gates, which also include the exactly $n$ outputs of gates that connect all the way to the end of the circuit. As a result, the number of internal links within the circuit is given by $N_{\ell}=3\times{\cal K}(P)-n$. For each of these internal links one can negate the output of the gate that acts before and the input of the gate that acts after, thus obtaining different gates that are part of a circuit that yields the same $P$. These negations can be implemented or not for each of the internal links, which leads to a degeneracy of at least $2^{N_{\ell}}$. For the polynomially-sized programs of interest in this paper, ${\cal K}(P)\geq n$, and we thus obtain
$S(P,{\cal K}(P))\geq 3{\cal K}(P)-n\geq 2{\cal K}(P)\;.$ (29)
Together with upper bound of Eq. (5) in the main text, the above lower bound establishes the extensively of $S(P,{\cal K}(P))$ with ${\cal K}(P)$, which we express generically as,
${\cal S}(P,{\cal K}(P))\approx\gamma(P;G)\;{\cal K}(P)\;,$ (30)
where the dependence on the specific permutation $P$ and the gate set $G$ are made explicit in the coefficient $\gamma(P;G)$. In writing this expression we tacitly assume that the argument given above will go through for any universal gate set G.
We note that the entropy computed above is associated with an internal redundancy that arises through a “gauge” transformation that is absorbed in redefining two gates that share a link. Beyond this minimum set of redundancies, which occurs for any arrangement of gates, there are other “structural” sources of entropy associated with different gate arrangements or circuit architectures that yield functionally equivalent circuits.
C - Model for the dependence of the circuit entropy on the depth and complexity of computation
@@
where $C(n,0)=\binom{n}{n/2}-\binom{n}{n/2+1}$ is a Catalan number that counts the number of balanced non-negative walks, i.e., walks that start and end at $k=0$ after $n$ (even) steps but that can only explore points with $k\geq 0$. By using Pascal’s triangle relations it is then easy to see that the solution to the recursion relation of  (32) is given by
@@
Let us consider the expression (35) with $p_{+}\gg p_{-},p_{0}$, a limit which is justified because for a generic circuit one expects that, upon addition a new (random) gate the complexity is much more likely to increase than to decrease or even remain unchanged. Also, in this limit $n_{+}\gg n_{-},n_{0}$ and $\frac{n_{+}-n_{-}+1}{n_{+}+1}\to 1$, leading to the following expression for the normalization,
@@
(42) then reduces to: $\bar{S}(0,{\cal N})\geq{\cal N}\;\log|G|-\frac{1}{2}\;{\cal N}\;\log|G|=\frac{1}{2}\;{\cal N}\;\log|G|$, which recovers the lower bound for the entropy of circuits implementing the identity permutation that we derived in the body of the paper by other means.